A company that makes clutch disks had the cash flows shown below for one department with a MARR of 25%. Calculate (a) the internal rate of return (Hint: 20%<=i*<=22%) and (b) the composite rate of return (Hint: 18%<=i’<=20%), using a reinvestment rate of 20% per year. Should the company continue to make clutch disks?
Year Cash flow, $
0 6000
1 -10000
2 12000
3 -10400
Answer #1

The given data is executed in an excel:

A 1 Year 3 Cash flow 6000 1 - 10000 2 12000 -10400 20.23% 19.95% 6 IRR 7 MIRR

a)

The IRR is 20.23% and

b)

The Composite rate of return is 19.95%

The company should not continue to make clutch disk as the IRR is Less than the given MARR of 25%

The excel formula and calculation used are shown below:
1 Year 20 31 4 2 5 3 6 IRR 7 MIRR Cash flow 6000 -10000 12000 - 10400 =IRR(B2:B5) =MIRR(B2:B5,20%,20%)

Please don't forget to rate the answer if its helpful, thank you.